# Quantum oscillation study of the large magnetoresistance in Mo substituted WTe2 single crystals

### v1

@@ III Results
Two of each Fermi-surface pocket is known to be present within the Brillouin zone, one on each side of the point [27, 28, 29]. As a result, the total electron concentration and the total hole concentration and their values are given in Table 3. The electron concentration is slightly higher than the hole concentration in the unsubstituted sample. However, with increasing Mo substitution, the oscillation frequencies change as described above and hence the hole concentration increases slightly while the electron concentration remains the same. This would imply that the charge compensation actually becomes better with Mo substitution.
@@ IV Discussion
The success of the two-band model lies in the fact that it predicts a parabolic and non-saturating MR () in case of perfect compensation (). However, the two-band model suffers from over parametrization and can lead to wrong conclusions as has been seen in the case of ZrSiSe where a Lifshitz transition inferred from the two-band fit has been later proved to be incorrect [60].
Considering this mutual dependency in the two-band model and the fact that it discards much information about the Fermi surface, it is clear that the measured changes in quantum oscillation frequencies give a more accurate picture of the evolution of the Fermi surface. A simple analysis of quantum oscillations suggests that the charge compensation improves with Mo substitution, at least up to . This last statement relies on the reliability of two assumptions underlying the estimate of the carrier densities: (i) that the prior band-structure calculations are correct in their assignment of the holes and electrons to the different Fermi-surface pockets, and (ii) that the pockets evolve isotropically in all cases, thus allowing the use of the spherical approximation. In any case, the decrease in MR that accompanies the increasing Mo substitution should be attributed solely or in part due to a decrease in transport mobility, as evidenced by the RRR data.
@@
In conclusion, we analyse the evolution of the quantum oscillations in WTe2 with Mo substitution. The MR decreases with substitution, but the charge compensation improves based on the quantum oscillation data. We attribute the MR decrease to the decrease in transport mobility, which is evident from the RRR values. Thus, we show that it is possible that the MR in WTe2 can decrease even with a very good level of charge compensation provided there is a fall in the transport mobility. Our quantum oscillation analysis further indicates that the quantum scattering time is unaffected by the Mo substitution.

### v3

@@ III Results
Two of each Fermi-surface pocket is known to be present within the Brillouin zone, one on each side of the point [27, 28, 29]. As a result, the total electron concentration and the total hole concentration and their values are given in Table 3. The electron concentration is slightly higher than the hole concentration in the unsubstituted sample. However, with increasing Mo substitution, the oscillation frequencies change as described above and hence the hole concentration increases slightly while the electron concentration remains the same. This would imply that the charge compensation actually becomes better with Mo substitution. The slightly different effect of the Mo substitution on the electron and hole pockets is likely due to different curvatures of the bands.
@@ IV Discussion
The success of the two-band model lies in the fact that it predicts a parabolic and non-saturating MR () in case of perfect compensation (). In this model, when the charge compensation is imperfect, the MR saturates at high fields, the value of which depends on the mobility. In fact, a very large mobility is counterproductive to a large MR, as it will lead to saturation at low field. For the optimum value of the mobility, the field at which the MR saturates can be beyond that accessible experimentally. However, the two-band model suffers from over parametrization and can lead to wrong conclusions as has been seen in the case of ZrSiSe where a Lifshitz transition inferred from the two-band fit has been later proved to be incorrect [60].
Considering this mutual dependency in the two-band model and the fact that it discards much information about the Fermi surface, it is clear that the measured changes in quantum oscillation frequencies give a more accurate picture of the evolution of the Fermi surface. A simple analysis of quantum oscillations suggests that the charge compensation improves with Mo substitution, at least up to . This last statement relies on the reliability of two assumptions underlying the estimate of the carrier densities: (i) that the prior band-structure calculations are correct in their assignment of the holes and electrons to the different Fermi-surface pockets, and (ii) that the pockets evolve isotropically in all cases, thus allowing the use of the spherical approximation. In any case, the decrease in MR that accompanies the increasing Mo substitution should be attributed solely or in part due to a decrease in transport mobility, as evidenced by the RRR data. It is possible that doping with Mo, while nominally isoelectronic, might affect the stability of other charged defects in the material, and could lead to changes in the Fermi energy and hence the charge compensation in addition to those arising from changes in the lattice parameters. Any such changes will be captured by the evolution of the quantum oscillation frequencies.
@@
In conclusion, we analyse the evolution of the quantum oscillations in WTe2 with Mo substitution. The MR decreases with substitution, but the charge compensation improves based on the quantum oscillation data. We attribute the MR decrease to the decrease in transport mobility, which is evident from the RRR values. Thus, we show that it is possible that the MR in WTe2 can decrease even with a very good level of charge compensation provided there is a fall in the transport mobility. Our quantum oscillation analysis further indicates that the quantum scattering time is unaffected by the Mo substitution. Ref 42 reports changes in mobility in WTe2 nanoflakes on Mo doping, similar to our results. However, they also use the two-band model to conclude that the charge balance remains undisturbed. Our data on bulk crystals shows that this approach is not valid for this material.